OpenStudy (anonymous):

what is the interval set? |x-2|>=4

OpenStudy (anonymous):

I know that [6,\[\infty\]is one

OpenStudy (anonymous):

\[[6,\infty)\]

OpenStudy (anonymous):

and I am hoping the other is \[(-\infty,-6] is the other\]

OpenStudy (anonymous):

\[|x-2| \ge 4 \implies -4 \ge x-2 \ge 4 \implies -2 \ge x \ge 6 \rightarrow x \in (-\infty,-2] \cup [6,\infty)\]

OpenStudy (anonymous):

I just got that as well, ty you so much

OpenStudy (anonymous):

No problem :P Sorry for the mis type the first time.

OpenStudy (anonymous):

all good